# On the Lipschitz Continuity of Set Aggregation Functions and Neural Networks for Sets

### v3

@@ B.3.4 The sum function is not Lipschitz continuous with respect to EMD
Therefore, the EMD between $X$ and $Y$ is equal to $1$. Then, we have that:
@@ B.3.5 The sum function is Lipschitz continuous with respect to the matching distance
which concludes the proof. The sum function is thus Lipschitz continuous with respect to the matching distance and the Lipschitz constant is equal to $1$.
@@ B.3.6 The sum function is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the sum function is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$ and $\mathbf{v}_{2}=(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)=\max_{i\in[2]}\min_{j\in[1]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{2}-\mathbf{u}_{1}\|=\sqrt{d}\epsilon$. We also have that:
@@ B.3.8 The max function is not Lipschitz continuous with respect to the matching distance
Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of $2$ and $1$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-c,-c,\ldots,-c)^{\top}$, and $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}\|=\sqrt{d}\epsilon$. Let also $\textbf{v}_{\text{max}}$ and $\textbf{u}_{\text{max}}$ denote the vectors that emerge after applying max pooling across all points of $X$ and $Y$, respectively. We also have that:
@@ B.4.2 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Suppose that the mean function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>3L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}\}$ be two multisets, consisting of $3$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-\frac{c}{2},-\frac{c}{2},\ldots,-\frac{c}{2})^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=\big(\frac{c}{2},\frac{c}{2},\ldots,\frac{c}{2}\big)^{\top}$, $\mathbf{v}_{3}=\big(\frac{c+\epsilon}{2},\frac{c+\epsilon}{2},\ldots,\frac{c+\epsilon}{2}\big)^{\top}$ and $\mathbf{u}_{3}=\big(-\frac{c+\epsilon}{2},-\frac{c+\epsilon}{2},\ldots,-\frac{c+\epsilon}{2}\big)^{\top}$ Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\max_{j\in[3]}\min_{i\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\|\mathbf{v}_{1}-\mathbf{u}_{3}\|=\frac{\sqrt{d}\epsilon}{2}$. We also have that:
@@ B.4.4 The sum function is not Lipschitz continuous with respect to the Hausdorff distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Suppose that the sum function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}\}$ be two multisets, consisting of $3$ vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-\frac{c}{2},-\frac{c}{2},\ldots,-\frac{c}{2})^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=\big(\frac{c}{2},\frac{c}{2},\ldots,\frac{c}{2}\big)^{\top}$, $\mathbf{v}_{3}=\big(\frac{c+\epsilon}{2},\frac{c+\epsilon}{2},\ldots,\frac{c+\epsilon}{2}\big)^{\top}$ and $\mathbf{u}_{3}=\big(-\frac{c+\epsilon}{2},-\frac{c+\epsilon}{2},\ldots,-\frac{c+\epsilon}{2}\big)^{\top}$ Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\max_{j\in[3]}\min_{i\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\|\mathbf{v}_{1}-\mathbf{u}_{3}\|=\frac{\sqrt{d}\epsilon}{2}$. We also have that:
@@ B.4.5 The max function is Lipschitz continuous with respect to the matching distance
The max function restricted to inputs from set $\mathcal{X}$ is thus Lipschitz continuous with respect to the matching distance and the Lipschitz constant is equal to $1$.
@@ B.5.1 The att mechanism is not Lipschitz continuous with respect to EMD
Suppose that the att mechanism is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(d\epsilon))L\epsilon}{\exp(d\epsilon)-1}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of $2$ vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{EMD}(X,Y)=\frac{1}{2}\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We define $g$ as the ReLU function. The attention coefficients of the elements of $X$ are equal to:
@@ B.5.2 The att mechanism is not Lipschitz continuous with respect to the matching distance
Suppose that the att mechanism is Lipschitz continuous with respect to the matching distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(d\epsilon))2L\epsilon}{\exp(d\epsilon)-1}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of $2$ vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=2\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We choose $g$ to be the ReLU activation function. The attention coefficients of the elements of $X$ are equal to:
@@ B.5.3 The att mechanism is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the att mechanism is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(d\epsilon))2L\epsilon}{\exp(d\epsilon)-1}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of $2$ vectors of dimension $d$. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)\leq\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=2\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We choose $g$ to be the ReLU activation function. The attention coefficients of the elements of $X$ are equal to:
@@ B.6.1 The $\textsc{att}_{\ell_{2}}$ mechanism is not Lipschitz continuous with respect to EMD
Suppose that the $\textsc{att}_{\ell_{2}}$ mechanism is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(-\sqrt{d}\epsilon))L\epsilon}{1-\exp(-\sqrt{d}\epsilon)}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of $2$ vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{EMD}(X,Y)=\frac{1}{2}\big\|\mathbf{v}_{2}-\mathbf{u}_{2}\big\|=\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix. We define $g$ as the ReLU function. The attention coefficients of the elements of $X$ are equal to:
@@
Suppose that the $\textsc{att}_{\ell_{2}}$ mechanism is Lipschitz continuous with respect to the matching distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(-\sqrt{d}\epsilon))2L\epsilon}{1-\exp(-\sqrt{d}\epsilon)}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of $2$ vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=2\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix. We choose $g$ to be the ReLU activation function. The attention coefficients of the elements of $X$ are equal to:
@@
Suppose that the $\textsc{att}_{\ell_{2}}$ mechanism is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(-\sqrt{d}\epsilon))2L\epsilon}{1-\exp(-\sqrt{d}\epsilon)}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of $2$ vectors of dimension $d$. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)\leq\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=2\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We choose $g$ to be the ReLU activation function. The attention coefficients of the elements of $X$ are equal to:

### v4

@@ B.3.4 The sum function is not Lipschitz continuous with respect to EMD
Therefore, the EMD between $X$ and $Y$ is equal to 11. Then, we have that:
@@ B.3.5 The sum function is Lipschitz continuous with respect to the matching distance
which concludes the proof. The sum function is thus Lipschitz continuous with respect to the matching distance and the Lipschitz constant is equal to 11.
@@ B.3.6 The sum function is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the sum function is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of 22 and 11 vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$ and $\mathbf{v}_{2}=(c+\epsilon,c+\epsilon,\ldots,c+\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)=\max_{i\in[2]}\min_{j\in[1]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{2}-\mathbf{u}_{1}\|=\sqrt{d}\epsilon$. We also have that:
@@ B.3.8 The max function is not Lipschitz continuous with respect to the matching distance
Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1}\}$ be two multisets, consisting of 22 and 11 vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-c,-c,\ldots,-c)^{\top}$, and $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}\|=\sqrt{d}\epsilon$. Let also $\textbf{v}_{\text{max}}$ and $\textbf{u}_{\text{max}}$ denote the vectors that emerge after applying max pooling across all points of $X$ and $Y$, respectively. We also have that:
@@ B.4.2 The mean function is not Lipschitz continuous with respect to the Hausdorff distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Suppose that the mean function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>3L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}\}$ be two multisets, consisting of 33 vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-\frac{c}{2},-\frac{c}{2},\ldots,-\frac{c}{2})^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=\big(\frac{c}{2},\frac{c}{2},\ldots,\frac{c}{2}\big)^{\top}$, $\mathbf{v}_{3}=\big(\frac{c+\epsilon}{2},\frac{c+\epsilon}{2},\ldots,\frac{c+\epsilon}{2}\big)^{\top}$ and $\mathbf{u}_{3}=\big(-\frac{c+\epsilon}{2},-\frac{c+\epsilon}{2},\ldots,-\frac{c+\epsilon}{2}\big)^{\top}$ Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\max_{j\in[3]}\min_{i\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\|\mathbf{v}_{1}-\mathbf{u}_{3}\|=\frac{\sqrt{d}\epsilon}{2}$. We also have that:
@@ B.4.4 The sum function is not Lipschitz continuous with respect to the Hausdorff distance
Let $\mathcal{X}$ denote a set that contains multisets of vectors of equal cardinalities, i. e., $|X|=M$ and $\in\mathcal{S}_{M}(\mathbb{R}^{d}), X𝒳\forall$ $\in\mathcal{X}$ where $M\in\mathbb{N}$. Suppose that the sum function restricted to inputs from set $\mathcal{X}$ is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>L\epsilon$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2},\mathbf{u}_{3}\}$ be two multisets, consisting of 33 vectors of dimension $d$, respectively. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(-\frac{c}{2},-\frac{c}{2},\ldots,-\frac{c}{2})^{\top}$, $\mathbf{v}_{2}=\mathbf{u}_{2}=\big(\frac{c}{2},\frac{c}{2},\ldots,\frac{c}{2}\big)^{\top}$, $\mathbf{v}_{3}=\big(\frac{c+\epsilon}{2},\frac{c+\epsilon}{2},\ldots,\frac{c+\epsilon}{2}\big)^{\top}$ and $\mathbf{u}_{3}=\big(-\frac{c+\epsilon}{2},-\frac{c+\epsilon}{2},\ldots,-\frac{c+\epsilon}{2}\big)^{\top}$ Clearly, we have that $d_{H}(X,Y)=\max_{i\in[3]}\min_{j\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\max_{j\in[3]}\min_{i\in[3]}\|\mathbf{v}_{i}-\mathbf{u}_{j}\|=\|\mathbf{v}_{3}-\mathbf{u}_{2}\|=\|\mathbf{v}_{1}-\mathbf{u}_{3}\|=\frac{\sqrt{d}\epsilon}{2}$. We also have that:
@@ B.4.5 The max function is Lipschitz continuous with respect to the matching distance
The max function restricted to inputs from set $\mathcal{X}$ is thus Lipschitz continuous with respect to the matching distance and the Lipschitz constant is equal to 11.
@@ B.5.1 The att mechanism is not Lipschitz continuous with respect to EMD
Suppose that the att mechanism is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(d\epsilon))L\epsilon}{\exp(d\epsilon)-1}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of 22 vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{EMD}(X,Y)=\frac{1}{2}\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We define $g$ as the ReLU function. The attention coefficients of the elements of $X$ are equal to:
@@ B.5.2 The att mechanism is not Lipschitz continuous with respect to the matching distance
Suppose that the att mechanism is Lipschitz continuous with respect to the matching distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(d\epsilon))2L\epsilon}{\exp(d\epsilon)-1}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of 22 vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=2\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We choose $g$ to be the ReLU activation function. The attention coefficients of the elements of $X$ are equal to:
@@ B.5.3 The att mechanism is not Lipschitz continuous with respect to the Hausdorff distance
Suppose that the att mechanism is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(d\epsilon))2L\epsilon}{\exp(d\epsilon)-1}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of 22 vectors of dimension $d$. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)\leq\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=2\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We choose $g$ to be the ReLU activation function. The attention coefficients of the elements of $X$ are equal to:
@@ B.6.1 The $\textsc{att}_{\ell_{2}}$ mechanism is not Lipschitz continuous with respect to EMD
Suppose that the $\textsc{att}_{\ell_{2}}$ mechanism is Lipschitz continuous with respect to EMD. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(-\sqrt{d}\epsilon))L\epsilon}{1-\exp(-\sqrt{d}\epsilon)}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of 22 vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{EMD}(X,Y)=\frac{1}{2}\big\|\mathbf{v}_{2}-\mathbf{u}_{2}\big\|=\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix. We define $g$ as the ReLU function. The attention coefficients of the elements of $X$ are equal to:
@@
Suppose that the $\textsc{att}_{\ell_{2}}$ mechanism is Lipschitz continuous with respect to the matching distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(-\sqrt{d}\epsilon))2L\epsilon}{1-\exp(-\sqrt{d}\epsilon)}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of 22 vectors of dimension $d$. Then, suppose that $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{M}(X,Y)=\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=2\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix. We choose $g$ to be the ReLU activation function. The attention coefficients of the elements of $X$ are equal to:
@@
Suppose that the $\textsc{att}_{\ell_{2}}$ mechanism is Lipschitz continuous with respect to the Hausdorff distance. Let $L>0$ be given. Let also $\epsilon>0$ and $c>\frac{2(1+\exp(-\sqrt{d}\epsilon))2L\epsilon}{1-\exp(-\sqrt{d}\epsilon)}$. Let $X=\{\mathbf{v}_{1},\mathbf{v}_{2}\}$, $Y=\{\mathbf{u}_{1},\mathbf{u}_{2}\}$ be two multisets, each consisting of 22 vectors of dimension $d$. Then, we set $\mathbf{v}_{1}=\mathbf{u}_{1}=(c,c,\ldots,c)^{\top}$, $\mathbf{v}_{2}=(\epsilon,\epsilon,\ldots,\epsilon)^{\top}$ and $\mathbf{u}_{2}=(-\epsilon,-\epsilon,\ldots,-\epsilon)^{\top}$. Clearly, we have that $d_{H}(X,Y)\leq\|\mathbf{v}_{2}-\mathbf{u}_{2}\|=2\sqrt{d}\epsilon$. Let also $\mathbf{W}=-\mathbf{I}$ and $\mathbf{q}=\mathbf{1}$ where $\mathbf{I}$ denotes the $d\times d$ identity matrix and $\mathbf{1}$ denotes the $d$-dimensional vector of all ones. We choose $g$ to be the ReLU activation function. The attention coefficients of the elements of $X$ are equal to: